OpenStudy (anonymous):

How do you do this log_8(x+6)=1-log_8(x+4) It says -2 but i have no idea how to get there

OpenStudy (anonymous):

\[\log_{8}(x+6)=1-\log_{8}(x+4) \]

OpenStudy (anonymous):

how is that true, if something has the same base, then it is = meaning (x+6)=1-( X+4)

OpenStudy (anonymous):

if you have 3^X+2 and 9^X+2 then u can drop it to the same base and say it is x+2=2x+4

OpenStudy (anonymous):

how can u just put them on one side, its an equality

OpenStudy (anonymous):

yea i was pretty sure u were wrong there

OpenStudy (anonymous):

did eveyone give up

OpenStudy (anonymous):

\[log_8(x+6)=1-Log_8(x+4)\] \[log_8(x+6)+Log_8(x+4)=1\] \[Log_8(x^2+10x+24)=1\] to the 8 \[x^2+10x+24=8\] \[x^2+10x+16=\] (x+2)(x+8) x=-2,-8

OpenStudy (anonymous):

ok well im sitting here with a mth teacher who says both of u are wrong

OpenStudy (anonymous):

can't be -8 because that would make log of negative number so only -2 is right answer

OpenStudy (anonymous):

its wrong according to him, u did not take into account the negative infront of the log and i have no idea what any of u are talking about

OpenStudy (akshay_budhkar):

Yes imran is right.. by that your professor means you have to check the domain and see if the value in log is not negative

OpenStudy (anonymous):

I did that after I got my numbers

OpenStudy (anonymous):

no he says u are both wrong and both answers are wrong

OpenStudy (anonymous):

u cant just swap sides of an equality regardless if it has the same base

OpenStudy (akshay_budhkar):

What is that imran is wrong in <.< the answer is -2 and that is what he has got!

OpenStudy (anonymous):

the log is NEGATIVE

OpenStudy (akshay_budhkar):

it is not!! if you have -2=x

OpenStudy (anonymous):

ITs not -2 im sitting next to a phd right now u wanna tell him he is wrong?

OpenStudy (akshay_budhkar):

lol even i have phd :P jkjk

OpenStudy (anonymous):

If a phd is sitting next to you , why are you asking for help here

OpenStudy (anonymous):

the log is an equality u can not just swap sides, and even if u could how can u just switch signs from a negative to a postive it makes NO sense

OpenStudy (akshay_budhkar):

lol imran :P

OpenStudy (anonymous):

because he wont give me the answer if he would i wouldnt be here trust me i have better things to do on friday night

OpenStudy (akshay_budhkar):

switch signs from a negative to a postive... what do u mean by that o.O

OpenStudy (anonymous):

u just moved the =1 over and said the log is no longer negative,

OpenStudy (anonymous):

even wolfram alpha http://www.wolframalpha.com/input/?i=log_8%28x%2B6%29%3D1-log_8%28%28x%2B4%29%29

OpenStudy (akshay_budhkar):

i still dont get what you want to say man!! Resubsititute the value -2 and confrim if u dun believe us <.<

OpenStudy (anonymous):

im not saying i dont beleieve u at all, im just saying my idiot phd professor says we are all wrong, i dunno why, besides that fact that he is an idiot

OpenStudy (akshay_budhkar):

lol show him wolfram alphas answer

OpenStudy (akshay_budhkar):

So he agreed o.O

OpenStudy (anonymous):

nope awesome now he says we are all missing something grrr i hate this guy

OpenStudy (akshay_budhkar):

Even wolfram alpha is missing something!?

OpenStudy (anonymous):

accroding to this crackpot

OpenStudy (akshay_budhkar):

or else review ur question.. u might have given an incomplete question here

OpenStudy (anonymous):

whats what i thought but i just double checked and that is exactly what is in the book

OpenStudy (akshay_budhkar):

Well i guarantee it is right.. you see how u handle this professor.. and if he still doesnt agree he can post the reason here... according to what i know it is RIGHT

OpenStudy (anonymous):

its driving me nuts, sigh as soon as he cracks and gives me the answer ill let u guys know